# Curves on Hirzebruch surfaces and semistability

### v2

@@ 1. Introduction
Let $(C,p_{1}\ldots,p_{n})$ be a general, pointed curve of genus $g$, by which we mean a general point of ${\mathcal{M}}_{g,n}$. Let $X$ be a smooth, projective variety. Can $C$ be interpolated through $n$ general points of $X$? If so, then in how many ways?
@@
Suppose that every irreducible component $Z\subset{\mathcal{M}}_{g,n}(X,\beta)$ dominating ${\mathcal{M}}_{g,n}\times X^{n}$ is generically smooth of expected dimension. Then, the geometric Tevelev degree ${\mathsf{Tev}}^{X}_{g,n,\beta}$ is by definition the degree of $\tau$.
@@
We also say that ${\mathsf{Tev}}^{X}_{g,n,\beta}$ is well-defined if the hypothesis of Definition 1.1 holds. Given a dominating component $Z\subset{\mathcal{M}}_{g,n}(X,\beta)$ that is generically smooth of expected dimension, a general curve interpolates through $n$ general points in $X$ via a map corresponding to a point of $Z$. Furthermore, this map admits no first-order deformations fixing $(C,p_{1},\ldots,p_{n})$ and the given points on $X$. If all dominating components $Z$ are generically smooth of expected dimension, then the total number of interpolating maps is given by ${\mathsf{Tev}}^{X}_{g,n,\beta}$.
@@ 1.2. Acknowledgments
We thank Roya Beheshti, Gavril Farkas, Eric Larson, Brian Lehmann, Aitor Iribar López, Rahul Pandharipande, Eric Riedl, Jason Starr, Sho Tanimoto, Richard Thomas, and Isabel Vogt for conversations related to this note. A.C. is supported by SNF grant P500PT-222363. C.L. has been supported by NSF Postdoctoral Fellowship DMS-2001976 and an AMS-Simons travel grant.
@@ 2. Semistability and interpolation
Let $Z\subset{\mathcal{M}}_{g,n}(X,\beta)$ be a component dominating $X^{n}\times{\mathcal{M}}_{g.n}$. Assume that $Z$ is generically smooth of expected dimension, and that (1) holds. Then, for a general point $[f:C\to X]$ of $Z$, the restricted tangent bundle $f^{*}T_{X}$ is semistable.
@@
  • When $X$ is a hypersurface of degree at most $(r+4)/3$ in $\mathbb{P}^{r+1}$ and $\beta$ is sufficiently large, then virtual Tevelev degrees of $X$ are enumerative (and thus equal to the geometric degrees) [2, Theorem 1.12(2)], and non-zero [4, Theorem 1.4, Theorem 1.5]. Therefore, for any general point of a component of ${\mathcal{M}}_{g,n}(X,d)$ dominating $X^{n}\times{\mathcal{M}}_{g,n}$, the restricted tangent bundle $f^{*}T_{X}$ is semistable for $f:C\to X$. While tangent bundles of smooth hypersurfaces are known to be semistable [23, Theorem 2], the stronger statement for restricted tangent bundles seems not previously to have been known.
@@ 3. Hirzebruch surfaces
Suppose further that $n\geq 2$ (or that $g=0$ and $n\geq 3$). Then, the component $Z$ is generically smooth of expected dimension.
@@
Note that $\beta$ necessarily has strictly positive anti-canonical degree on ${\mathcal{H}}_{a}$. By standard results on Severi varieties of surfaces (see e.g. [9, Proposition 2.4]), a component $Z$ can only fail to be generically smooth of expected dimension if every point of $Z$ parametrizes a map $f$ whose image has anti-canonical degree at most 3. The only possible images of such maps are the fibers of $\pi:{\mathcal{H}}_{a}\to\mathbb{P}^{1}$ and the infinity-section. However, this is impossible if $f$ is required to interpolate through 2 or more general points of ${\mathcal{H}}_{a}$. $\Diamond$
@@
As $C$ is general, the space ${\mathcal{M}}_{m}(C,\mathbb{P}^{1})$ is pure of expected dimension, and ev has relative dimension 0. Thus, the locus on ${\mathcal{M}}_{m}(C,\mathbb{P}^{1})\times C^{n}$ where the conclusion of Lemma 3.4 cannot dominate the target. $\Diamond$

### v3

@@ 1. Introduction
Let ${\mathcal{M}}_{g,n}$ be the Deligne-Mumford moduli space of genus $g$, $n$-pointed curves. Let $(C,p_{1}\ldots,p_{n})$ be a general, pointed curve of genus $g$, by which we mean a general point of ${\mathcal{M}}_{g,n}$. Let $X$ be a smooth, projective variety. Can $C$ be interpolated through $n$ general points of $X$? If so, then in how many ways?
@@
Suppose that every irreducible component $Z\subset{\mathcal{M}}_{g,n}(X,\beta)$ dominating ${\mathcal{M}}_{g,n}\times X^{n}$ is generically smooth of the expected dimension. Then, the geometric Tevelev degree ${\mathsf{Tev}}^{X}_{g,n,\beta}$ is by definition the degree of $\tau$.
@@
We also say that ${\mathsf{Tev}}^{X}_{g,n,\beta}$ is well-defined if the hypothesis of Definition 1.1 holds. Given a dominating component $Z\subset{\mathcal{M}}_{g,n}(X,\beta)$ that is generically smooth of the expected dimension, a general curve interpolates through $n$ general points in $X$ via a map corresponding to a point of $Z$. Furthermore, this map admits no first-order deformations fixing $(C,p_{1},\ldots,p_{n})$ and the given points on $X$. If all dominating components $Z$ are generically smooth of the expected dimension, then the total number of interpolating maps is given by ${\mathsf{Tev}}^{X}_{g,n,\beta}$.
@@ 1.2. Acknowledgments
We thank Roya Beheshti, Gavril Farkas, Eric Larson, Brian Lehmann, Aitor Iribar López, Rahul Pandharipande, Eric Riedl, Jason Starr, Sho Tanimoto, Richard Thomas, and Isabel Vogt for conversations related to this note, and the anonymous referee for their helpful comments. A.C. is supported by SNF grant P500PT-222363. C.L. has been supported by NSF Postdoctoral Fellowship DMS-2001976 and an AMS-Simons travel grant.
@@ 2. Semistability and interpolation
Let $Z\subset{\mathcal{M}}_{g,n}(X,\beta)$ be a component dominating ${\mathcal{M}}_{g.n}\times X^{n}$. Assume that $Z$ is generically smooth of the expected dimension, and that (1) holds. Then, for a general point $[f:C\to X]$ of $Z$, the restricted tangent bundle $f^{*}T_{X}$ is semistable.
@@
  • When $X$ is a hypersurface of degree at most $(r+4)/3$ in $\mathbb{P}^{r+1}$ and $\beta$ is sufficiently large, then virtual Tevelev degrees of $X$ are enumerative (and thus equal to the geometric degrees) [2, Theorem 1.12(2)], and non-zero [4, Theorem 1.4, Theorem 1.5]. Therefore, for any general point of a component of ${\mathcal{M}}_{g,n}(X,d)$ dominating ${\mathcal{M}}_{g,n}\times X^{n}$, the restricted tangent bundle $f^{*}T_{X}$ is semistable for $f:C\to X$. While tangent bundles of smooth hypersurfaces are known to be semistable [23, Theorem 2], the stronger statement for restricted tangent bundles seems not previously to have been known.
@@ 3. Hirzebruch surfaces
Suppose further that $n\geq 2$ (or that $g=0$ and $n\geq 3$). Then, the component $Z$ is generically smooth of the expected dimension.
@@
Note that $\beta$ necessarily has strictly positive anti-canonical degree on ${\mathcal{H}}_{a}$. By standard results on Severi varieties of surfaces (see e.g. [9, Proposition 2.4]), a component $Z$ can only fail to be generically smooth of the expected dimension if every point of $Z$ parametrizes a map $f$ whose image has anti-canonical degree at most 3. The only possible images of such maps are the fibers of $\pi:{\mathcal{H}}_{a}\to\mathbb{P}^{1}$ and the infinity-section. However, this is impossible if $f$ is required to interpolate through 2 or more general points of ${\mathcal{H}}_{a}$. $\Diamond$
@@
As $C$ is general, the space ${\mathcal{M}}_{m}(C,\mathbb{P}^{1})$ is pure of the expected dimension, and ev has relative dimension 0. Thus, the locus on ${\mathcal{M}}_{m}(C,\mathbb{P}^{1})\times C^{n}$ where the conclusion of Lemma 3.4 fails cannot dominate the target. $\Diamond$